OpenStudy (anonymous):

could someone please help me with these two.. ): i'm studying for a test and i got stuck with these too.. i don't want to skip it because it might be on the test. solve. \[4^{5x} = 16 ^{2x-1}\] and \[3^{5x}9^{x ^{2}} = 27\]

OpenStudy (anonymous):

Hint: 16 = 4²

OpenStudy (anonymous):

27 = 3³

OpenStudy (anonymous):

@alexeis_nicole smart move! it's highly chance to be on the test :)

OpenStudy (anonymous):

ohh. okay so for the first one... would it be \[4^{5x} = 4^{2(2x-1)}\]\[4^{5x} = 4^{4x-2}\]\[5x=4x-2\]\[5x-4x=-2\]\[x=-2\] right??? :)

OpenStudy (anonymous):

Perfect :)

OpenStudy (anonymous):

yay ! k i'll try the next one then get back to you (x

OpenStudy (anonymous):

k.. uh... what do i do with the \[x^2\] in the second equation ?? :S

OpenStudy (anonymous):

9 = 3²

OpenStudy (anonymous):

soo... am i supposed to get 5x + 2x^2 = 3 ? :S

OpenStudy (anonymous):

Yes, can you solve 2x² + 5x - 3 = 0

OpenStudy (anonymous):

uhmmm.. id just factor this right? sorry,, i just got lost Dx

OpenStudy (anonymous):

5x + 2x² = 3 2x² + 5x - 3 = 0 Can you apply quadratic formula?

OpenStudy (anonymous):

i applied the quadratic formula and i got 3, and -0.5

OpenStudy (anonymous):

oH WAIT... i did something wrong.. let me try that again )x sorry!

OpenStudy (anonymous):

i got 0.5 and -3.

OpenStudy (anonymous):

Correct now :)

OpenStudy (anonymous):

so... would that be the answer? :S

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

hmm... :/ the book tells me that its supposed to be -.02 ?

OpenStudy (anonymous):

If so, check your post?

OpenStudy (anonymous):

hmm... its the right equation... /:

OpenStudy (anonymous):

If your post correct, then the solution has typo! Point it out to your teacher :)

OpenStudy (anonymous):

hmm probably. yeah i will, thank you :)

OpenStudy (anonymous):

Be confident ;)